OpenStudy (anonymous):

Solve for x: 3 cos x= 9cosxtan^2x; -∞≤0≤ ∞

OpenStudy (loser66):

Question: why do we have cos x at both sides? if it's the right problem, can we cancel them out?

OpenStudy (anonymous):

I think so I don't really know

OpenStudy (loser66):

so, neither I. @Luigi0210 can you help us?

OpenStudy (luigi0210):

Have you tried trig identities?

OpenStudy (anonymous):

I couldn't figure out which one to use

OpenStudy (luigi0210):

tan^2x=sec^2x-1

OpenStudy (luigi0210):

Do you have any ideas @Loser66 ?

OpenStudy (loser66):

I have a weird idea like this \[3cos x = 9 cos x tan^2 x\\\frac{1}{3}= tan^2x=\frac{sin^2x}{cos^2x}\\3sin^2x= cos^2x\\3sin^2x -cos^2x=0\\3-3cos^2x-cos^2x=0\\3-4cos^2x=0\\cos^2x = \frac{3}{4}\]and then I am stuck, hehehehe...

OpenStudy (loser66):

@Jazzyone2 can you see my mistake from the stuff?

OpenStudy (luigi0210):

Yea! I had the same idea after the identity.

OpenStudy (anonymous):

Yeah i got it